If an iron wire is drawn to decrease it's radius by 0.5 precent of increase in its resistance will be
1
Expert's answer
2019-07-19T14:37:37-0400

The resistance of a wire of length "l" and cross-sectional area "A" is given by

"R=\rho \frac{l}{A}"

where "\rho" is the resistivity of the material of the wire.

If radius is decreased by 0.5 % then new radius "r'=0.995r"

New area "A'=\pi r'^2=(0.995)^2\pi r^2 = (0.995)^2A"

New resistance "R'=\rho \frac{l}{(0.995)^2A}=\frac{R}{0.995^2}=1.01R"

Percentage increase in resistance

"=\frac{R'-R}{R}\times 100\\
=\frac{1.01R-R}{R}\times 100\\
=1"

Increase in resistance = 1 %